ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 17, 46 Usmanov, Shukhrat M., Academy of Sciences of Uzbekistan, Tashkent, Uzbekistan Operator valued weights on real von Neumann algebras and JW - algebras Operator valued weights on real von Neumann algebras and reversible JW - algebras are defined. Properties of its existence and uniqueness are announced. New formulate of theorem on existence of conditional expectations on the reversible JW -algebras are obtained.